D | Question 15 5 pts A taxicab moves five blocks due north, five blocks due east, and another three blocks due north. Assume all blocks are of equal size. What is the magnitude of the taxis displacement, start to finish? O 12 blocks 。8.6 blocks O 9.8 blocks O 9.4 blocks
Answer #1

First displacement, d1 = 5j (North)
Second displacement, d2 = 5i (East)
Third displacement, d3 = 3j (North)

Total displacement = d1+d2+d3 = 5i+8j
Magnitude = sqrt(5^2+8^2) = sqrt(25+64) = sqrt (89) = 9.4 blocks
Answer is 9.4 blocks